2/(x+2) +2 = 7/ (x+2)
1
Expert's answer
2014-07-23T11:22:15-0400
Subtract 2/(x+2) from both parts and obtain 2=7/(x+2)-2/(x+2), collect similar terms, 2=5/(x+2). Multiply by (x+2), which is non-zero and obtain 2(x+2)=5. Dividing by 2 results in x+2=5/2. Next, subtract 2 from both sides x=5/2-2.
Answer: x=1/2.